OpenStudy (anonymous):

Evaluate the limit as x approaches -4. ((1/4)+(1/x))/(4+x)

OpenStudy (anonymous):

\[1/4+1/x=(x+4)/4x\]

OpenStudy (turingtest):

\[\lim_{x\to-4}{\frac14+\frac1x\over4+x}\]simplify the numerator

OpenStudy (anonymous):

(x+4/4x)/(4+x)?

OpenStudy (anonymous):

then take out -1?

OpenStudy (anonymous):

meaning x+4 goes out

OpenStudy (anonymous):

what you mean x+4 goes out?

OpenStudy (anonymous):

because you are dividing by 4+x which is the same as x+4 on top \[(x+4)/4x/(4+x)=(x+4)/4x*1/(x+4)=1/4x\]

OpenStudy (anonymous):

so it becomes lim x-> -4 = 1/4x then substitute?

OpenStudy (anonymous):

ofcourse

OpenStudy (anonymous):

-1/16?

OpenStudy (anonymous):

great

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

never mind theres no problem its correct

OpenStudy (anonymous):

are you sure...?

OpenStudy (anonymous):

for sure